OpenStudy (anonymous):

Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = -226. (2 points)

OpenStudy (anonymous):

-334 -274 -284 -346

OpenStudy (haseeb96):

a = 26 a+ 21d=-226 26 +21d=-226 21d = -226 -26 21d = -252 d = -252/21 d= -12 now look at this one a+ 30d 26 +30(-12) =-334 is the correct answer

OpenStudy (jiteshmeghwal9):

\[a_1=a=26\]\[a_{22}=a+21d=-226\]\[a_{22}=26+21d=-226\] \[d=-252/21\]\[d=-12\]\[a_{31}=a+30d=26+30(-12)=?\]

OpenStudy (anonymous):

thanks so much both

OpenStudy (haseeb96):

haha I hope she will learn now.

OpenStudy (anonymous):

ill give you a medal this time since you answered last time and didnt get one

OpenStudy (haseeb96):

yes she is correct